# Measure-Theoretic Time-Delay Embedding

### v1

@@ Appendix A Proofs from Main Text
where $Y_{f(x)}=df_{x}X_{x}$ is a vector field on $N$. Moreover, since $d\varphi_{f(x)}$ lies in the cotangent bundle on $N$, we have $d\varphi_{y}Y_{y}=q_{y}(\nabla\varphi_{y},Y_{y})$. By plugging this into the equation above with $y=f(x)$, we obtain (9). ∎
@@
We proceed in two steps. First, let $\rho_{t}$ be an AC curve, and $v_{t}=\frac{d}{dt}\rho_{t}$ whose existence is guaranteed by Proposition 3. We then prove that $(F(\rho_{t}),\widetilde{dF}_{\rho_{t}}(v_{t}))$ satisfies the continuity equation and that $\int_{0}^{1}\|\widetilde{dF}_{\rho_{t}}(v_{t})\|_{L^{2}(F(\rho_{t}))}{\mathrm{d}}t<\infty$, which guarantees that $F$ is AC by Proposition 3. Secondly, we extend the previous results to the projected version $dF_{\rho}=P^{F(\rho)}\widetilde{dF_{\rho}}(v)$ which lies in $T_{F(\rho)}\mathcal{P}(N)$ by the definition of the projection operator ((8)).
@@
$\int_{M}q_{f(x)}(df_{x}(v_{t}(x)),df_{x}(v_{t}(x)))\,{\mathrm{d}}\rho_{t}(x)\leq\int_{M}\|df_{x}\|^{2}g_{x}(v_{t}(x),v_{t}(x))\,{\mathrm{d}}\rho_{t}(x)\leq\sup_{x\in M}\|df_{x}\|^{2}\int_{M}g_{x}(v_{t}(x),v_{t}(x))\,{\mathrm{d}}\rho_{t}(x)\,.$
Next, we show that $(F(\rho_{t}),\widetilde{dF}_{\rho_{t}}(v_{t}))$ satisfies the continuity equation ((5)). For any test function $\varphi\in\mathcal{D}(N\times[0,T])$,
@@
Since $\text{supp}(\varphi)$ is compact and $(f\otimes\text{{Id}})$ is proper, $(f\otimes\text{{Id}})^{-1}\left(\text{supp}(\varphi)\right)$ is also compact. Moreover, supp$(\widetilde{\varphi})$ is a closed subset of a compact set, and hence itself compact. Thus, $\widetilde{\varphi}\in\mathcal{D}(M\times[0,T])$. Since $(\rho_{t},v_{t})$ satisfies (4) by assumption, (S4) is always zero, which implies that $(F(\rho_{t}),\widetilde{dF}_{\rho_{t}}(v_{t}))$ satisfies the continuity equation ((5)). Finally, we turn to $dF_{\rho_{t}}(v_{t})$.
@@ B.1 Synthetic Examples
which are standard values known to produce chaotic trajectories. In Section 4.2 of the main text, we compared the pointwise ((13)) and measure-based ((14)) reconstruction schemes for sparse and noisy data coming from these systems.
Here, we show another numerical example, but this time, a large amount of clean, noise-free data is used instead. Figure 1 shows the results in which we perform the full state reconstruction of the same three systems. For this test, we partition the delay state into $100$ cells, which results in $100$ different probability measures for the measure-based loss ((14)). It is worth noting that each measure is an empirical distribution based on thousands of samples. To reduce the computational cost, we used mini-batching to reduce the number of samples in each of the $100$ empirical distributions from $5000$ to $50$. We remark that mini-batch training is not performed in the tests shown in the main text since all examples there are on sparse datasets where the problem instead becomes a lack of data.
After subdividing the training data, we then parameterize $\mathcal{R}_{\theta}$ as a two-layer feed-forward neural network with $100$ nodes in each layer and a hyperbolic tangent activation function. We enforce the distributional loss (14) during training. We use a learning rate of $10^{-3}$. Despite using only $K=100$ measures for learning $\mathcal{R}_{\theta}$, after $100$ epochs of training, we find that the network forecasts the full state of each dynamical system with high-precision. See Figure 5(a) for the Lorenz-63 system, Figure 5(b) for the Rössler system and Figure 5(c) for the Lotka–Volterra system.
Figure 5: Learning the full-state reconstruction map for three low-dimensional chaotic attractors. The measures $\Phi\#\mu_{i}$ on the delayed attractor are shown in the left column with each measure colored differently. On the right column, the neural network forecast during testing is shown as the dotted-black line, whereas the ground truth is in red. In this case, the data is noise-free.
@@ B.2 NOAA SST Reconstruction
In Section 4.3, we performed full state reconstruction on the NOAA SST dataset using both the pointwise and measure-based approaches. Figure 6 visualizes the dataset we use to perform the reconstruction. Figure 6(a) shows a single snapshot of the dataset, as well as the geospatial location at which we collect partial observational data. Figure 6(b) plots the temperature time series at this location. Figures 6(c) and 6(d) visualize projections of the input-output measure data used to train the model according to (14). More specifically, Figure 6(d) plots the input measures constructed according to the partially observed data in time-delay coordinates, while 6(c) shows the output measures in the POD reconstruction space.
@@ B.3 ERA5 Wind Field Reconstruction
In Section 4.4, we performed full state reconstruction on a portion of the ERA5 wind field dataset. Figure 7 visualizes a single snapshot of the dataset, including the randomly sampled geospatial locations at which we collect partial observational data, as well as an example wind speed time-series at one of these locations. Figure 8 visualizes results for performing the full-state reconstruction on the dataset using both the pointwise ((13)) and measure-based ((14)) approaches.

### v2

@@ Appendix A Proofs from the Main Text
where $Y_{f(x)}:=df_{x}X_{x}$ is a vector field on $N$. Moreover, since $d\varphi_{f(x)}$ lies in the cotangent bundle on $N$, we have $d\varphi_{y}Y_{y}=q_{y}(\nabla\varphi_{y},Y_{y})$. By plugging this into the equation above with $y=f(x)$, we obtain (9). ∎
@@
We proceed in two steps. First, let $\rho_{t}$ be an AC curve, and $v_{t}=\frac{d}{dt}\rho_{t}$ whose existence is guaranteed by Proposition 6. We then prove that $(F(\rho_{t}),\widetilde{dF}_{\rho_{t}}(v_{t}))$ satisfies the continuity equation and that $\int_{0}^{1}\|\widetilde{dF}_{\rho_{t}}(v_{t})\|_{L^{2}(F(\rho_{t}))}{\mathrm{d}}t<\infty$, which guarantees that $F$ is AC by Proposition 6. Secondly, we extend the previous results to the projected version $dF_{\rho}=P^{F(\rho)}\widetilde{dF_{\rho}}(v)$ which lies in $T_{F(\rho)}\mathcal{P}(N)$ by the definition of the projection operator (see (8)).
@@
$\int_{M}q_{f(x)}(df_{x}(v_{t}(x)),df_{x}(v_{t}(x)))\,{\mathrm{d}}\rho_{t}(x)$ $\leq\int_{M}\|df_{x}\|^{2}g_{x}(v_{t}(x),v_{t}(x))\,{\mathrm{d}}\rho_{t}(x)$
$\leq\sup_{x\in M}\|df_{x}\|^{2}\int_{M}g_{x}(v_{t}(x),v_{t}(x))\,{\mathrm{d}}\rho_{t}(x)\,.$
Next, we show that $(F(\rho_{t}),\widetilde{dF}_{\rho_{t}}(v_{t}))$ satisfies the continuity equation (5). For any test function $\varphi\in\mathcal{D}(N\times[0,T])$,
@@
[…]
@@ B.1 Synthetic Examples
which are standard values known to produce chaotic trajectories. In Section 4.2 of the main text, we compared the pointwise (see (14)) and measure-based (see (15)) reconstruction schemes for sparse and noisy data coming from these systems.
Here, we show another numerical example, but this time, a large amount of clean, noise-free data is used instead. Fig. 1 shows the results in which we perform the full state reconstruction of the same three systems. For this test, we partition the delay state into $100$ cells, which results in $100$ different probability measures for the measure-based loss (see (15)). It is worth noting that each measure is an empirical distribution based on thousands of samples. To reduce the computational cost, we used mini-batching to reduce the number of samples in each of the $100$ empirical distributions from $5000$ to $50$. We remark that mini-batch training is not performed in the tests shown in the main text since all examples there are on sparse datasets where the problem instead becomes a lack of data.
After subdividing the training data, we then parameterize $\mathcal{R}_{\theta}$ as a two-layer feed-forward neural network with $100$ nodes in each layer and a hyperbolic tangent activation function. We enforce the distributional loss (15) during training. We use a learning rate of $10^{-3}$. Despite using only $K=100$ measures for learning $\mathcal{R}_{\theta}$, after $100$ epochs of training, we find that the network reconstructs the full state of each dynamical system with high-precision. See Fig. 6(a) for the Lorenz-63 system, Fig. 6(b) for the Rössler system and Fig. 6(c) for the Lotka–Volterra system.
Figure 6: Learning the full-state reconstruction map for three low-dimensional chaotic attractors. The measures $\Phi\#\mu_{i}$ on the delayed attractor are shown in the left column with each measure colored differently. On the right column, the neural network reconstruction during testing is shown as the dotted-black line, whereas the ground truth is in red. In this case, the data is noise-free.
@@ B.2 NOAA SST Reconstruction
In Section 4.3, we performed full state reconstruction on the NOAA SST dataset using both the pointwise and measure-based approaches. Fig. 7 visualizes the dataset we use to perform the reconstruction. Fig. 7(a) shows a single snapshot of the dataset, as well as the geospatial location at which we collect partial observational data. Fig. 7(b) plots the temperature time series at this location. Figures 7(c) and 7(d) visualize projections of the input-output measure data used to train the model according to (15). More specifically, Fig. 7(d) plots the input measures constructed according to the partially observed data in time-delay coordinates, while 7(c) shows the output measures in the POD reconstruction space.
@@ B.3 ERA5 Wind Field Reconstruction
In Section 4.4, we performed full state reconstruction on a portion of the ERA5 wind field dataset. Fig. 8 visualizes a single snapshot of the dataset, including the randomly sampled geospatial locations at which we collect partial observational data, as well as an example wind speed time-series at one of these locations. Fig. 9 visualizes results for performing the full-state reconstruction on the dataset using both the pointwise (see (14)) and measure-based (see (15)) approaches.